OpenStudy (anonymous):

Having trouble solving a equation. 0.6w-0.4=-6+w

OpenStudy (anonymous):

hi there, i will go over the steps on how to solve this problem.

OpenStudy (anonymous):

First, add 0.4 to both sides. this should look like this: 0.6w-0.4=-6+w +0.4 +0.4 --------------- 0.62=-5.6+w

OpenStudy (anonymous):

can you solve from here?

OpenStudy (anonymous):

This is what I did 0.6w-0.4=-6+w -0.8w-0.4=-6+0 -0.8w-0.4=-6+0.4 than I divided

OpenStudy (anonymous):

and I got the answer 7

OpenStudy (anonymous):

It says the answer is 14 but I dont see how

OpenStudy (anonymous):

hmm..well let's see here. so far we have 0.62w=-5.6+w so subtract w from both sides 0.62w=-5.6+w -w -w --------------- -0.382w=-5.6

OpenStudy (anonymous):

do you see what i did here? I think you didn't add up your similar variables correctly. If I'm not mistaken

OpenStudy (anonymous):

if you divide these values, the answer is indeed 14 It's actually 14.357

OpenStudy (anonymous):

I don't think I understand where the -0.382 came from...I'm sorry something so simple and I'm making it hard really ridiculous that I can't get this done already

OpenStudy (anonymous):

lool don't worry about it! the -0.382 came from when you subtract 0.62w with -1w, or -w

OpenStudy (anonymous):

think about it this way: 0.62-1

OpenStudy (anonymous):

does that make sense? @Yogobro298

OpenStudy (anonymous):

yes it makes more sense now

OpenStudy (anonymous):

awesome! Glad to have helped :)